# Prediction Intervals for Interim Events in Randomized Clinical Trials with Time-to-Event Endpoints

### v1

@@ Abstract
Time-to-event endpoints are central to evaluate treatment efficacy across many disease areas. Many trial protocols include interim analyses within group-sequential designs that control type I error via spending functions or boundary methods. The corresponding operating characteristics depend on the number of looks and the information accrued. Planning interim analyses with time-to-event endpoints is challenging because statistical information depends on the number of observed events. Ensuring adequate follow-up to accrue the required events is therefore critical, making interim prediction of information at scheduled looks and at the final analysis essential. While several methods have been developed to predict the calendar time required to reach a target number of events, to the best of our knowledge there is no established framework that addresses the prediction of the number of events at a future date with corresponding prediction intervals. Starting from an prediction interval approach originally developed in reliability engineering for the number of future component failures, we reformulated and extended it to the context of interim monitoring in clinical trials. This adaptation yields a general framework for event-count prediction intervals in the clinical setting, taking the patient as the unit of analysis and accommodating a range of parametric survival models, patient-level covariates, stagged entry and possible dependence between entry dates and lost to follow-up. Prediction intervals are obtained in a frequentist framework from a bootstrap estimator of the predictive distribution of future events. The performance of the proposed approach is investigated via simulation studies and illustrated by analyzing a real-world phase III trial in childhood acute lymphoblastic leukaemia.
@@ 1 Introduction
Time-to-event endpoints are a necessary means for evaluating treatment efficacy in many disease areas. However, designing trials in which the primary endpoint is a time-to-event variable is challenging. In this context, the statistical information of interest depends on the number of events that occur rather than on the number of patients enrolled. Accordingly, the trial is usually powered on the expected number of events [80]. Interim analyses are increasingly planned to allow early stopping for efficacy or futility for ethical and economic reasons [77]. Group-sequential designs are typically adopted and type-I error is controlled via alpha-spending or boundary methods [42]. The operating characteristics of a group sequential test - power, error rates, and stopping probabilities - depend on the number and timing of looks, on the information accrued at each analysis and the maximum planned information at the final analysis. Consequently, the protocol must ensure sufficient follow‐up to achieve the target number of events required to preserve the nominal power. In addition, interim analyses should be conducted at the information levels specified at the design stage. If analyses are performed at times corresponding to a different number of events than planned, the Type I error probability may be adversely affected [46] [60]
@@
However, achieving both aims is difficult in practice. Sample size calculation relies on the correct specification of the survival distribution, yet trustworthy informations on survival estimates are often unavailable at the design stage and may shift as data mature. When the design assumptions prove inaccurate while the trial is ongoing, the study could jeopardize its operational characteristics. Because information accrues only when events occur, an unforeseen slowdown in event rates can shift interim analyses to information times that depart from the protocol plan or push completion well beyond the planned timeline. This has an impact on the budget and logistics of the trial, even when the nominal power of the trial is preserved. On the other hand, an unexpected increase in event rates may put the trial at risk of being over-run [78] and appropriate methods should be considered [7]. A disciplined interim monitoring framework can identify such discrepancies early and support timely corrective actions. However, because only a few interim looks are typically justifiable operating characteristics are usually evaluated under the originally planned interim scheme.
@@
In light of these challenges, it is important to prospectively predict while the trial is ongoing the information expected at each planned interim analysis, the maximum (final) information at the scheduled final analysis or the calendar time at which these targets will be reached. In a maximum-information design, the trial ends once the target number of events has accrued [44]. A natural predictand is therefore the calendar time required to reach that target, often termed milestone or time to endpoint maturation [57] [76]. In a maximum-duration design, follow-up is limited by a fixed calendar time [44]. In this case, natural predictands are the cumulative number of events or, equivalently, the additional number of events expected by future calendar dates and at the planned study end.
@@
The point and interval prediction of the time to endpoint maturation has been extensively studied in the literature [6, 83, 81, 82, 66, 16, 47, 52]. Heitjan et al. [34] provide a comprehensive review of existing methods, highlighting the importance of prediction intervals for time to endpoint maturation to quantify uncertainty around the study’s closing date. Some approaches accommodate ongoing accrual at the time of prediction [76] [85] [3]
@@
Predicting the additional number of events at interim time points is critical for decision-making, yet most available methods in the biostatistical literature provide only point prediction under limited survival distribution. Fang and Su [24] proposed a hybrid approach that combines nonparametric and parametric modeling to estimate the survival curve. Specifically, change points in the survival function are first detected using a data-driven algorithm developed by Goodman et al. [28]. The segment preceding the final change point is estimated nonparametrically, while the tail beyond it is modeled using a parametric exponential or Weibull distribution. This approach is implemented by Rufibach [63] in the eventTrack R package. A growing body of literature has also addressed the use of point prediction for the purpose of sample size re-estimation [79, 31, 72, 25, 56]. Even though point predictions are easy to communicate they do not quantify uncertainty about reaching the planned information level and target number of events by the prespecified date. Prediction intervals address this limitation by explicitly reflecting uncertainty through their limits and width, making them more informative for interim trial monitoring. To the best of our knowledge, only two approaches address directly the issue of providing prediction interval for the additional number of events. Anisimov [3] derived closed-form prediction intervals for the number of future events in ongoing multicenter time-to-event trials, explicitly accounting for both newly recruited and at-risk patients under varied recruitment scenarios. The approach models center-level accrual via a Poisson–gamma process and event dynamics with continuous-time finite Markov chains under an exponential model. Anisimov et al. [2] extended the framework to incorporate a cure fraction, under both exponential and Weibull assumptions. In this paper, we propose a general framework to construct prediction intervals for the additional number of events at interim analysis in fixed-duration clinical trials by adapting and extending a method originally developed in reliability engineering for predicting component failures [71]. We reformulate the approach for the clinical setting by treating the patient (rather than a component cohort) as the unit of analysis, accommodating possible dependence between entry dates and lost to follow-up and leveraging patient-level covariates available at unblinded interim looks, as in interim analyses for early futility or efficacy.
@@
This paper is organized as follows: Section 2 introduces the notation, the prediction framework and briefly reviews the methods originally proposed by Tian et al. [71] for obtain prediction intervals; Section 3 outlines our proposed adaptation and extensions; Section 4, presents the results of a simulation study to evaluate the performance of the modified framework; Section 5 include an application to a real-world phase III trial in childhood acute lymphoblast leukemia (ALL); Section 6 discusses the implications of our findings and highlights potential extensions.
@@ 2 Background
In this paper, we adapt and extend the direct bootstrap method in the multiple-chort scenario. The method has shown stronger empirical performance than the calibration bootstrap. Moreover, unlike the generalized pivotal-quantity bootstrap—whose construction is primarily well-defined under Type II censoring and only approximated under Type I censoring—the direct bootstrap can accommodates more general censoring mechanisms. We therefore focus on this approach and briefly describe it below.
@@ 2.1 Prediction Intervals Definition
Some comments on the relationship between $Y$ and $\bm{D}$ are in order. In new-sample prediction (also known as out-of-sample prediction), the predictand belongs to a future, independent sample under the assumed model, so $Y$ is independent of $\bm{D}$. In within-sample prediction (also known as forecast), the predictand represents a random variable belonging to the same sample for which the prediction is made in the future [55]. In this case, $Y$ and $\bm{D}$ are not independent. This scenario is common in both reliability analysis and interim monitoring settings.
@@ 2.2 The Direct-bootstrap, Multiple Chort Method (DB-MC)
At the time of the prediction $t_{c}$ there are $g$ groups of items of size $n_{j}$ put on test at different times, with j=1,gj=1...,g . In each group we observe $\delta_{j}=\sum_{i=1}^{n_{j}}\mathrm{I}(T_{i}\leq\tau_{j})$ failures and $\gamma_{j}=n_{j}-\delta_{j}$ items censored at $t_{c}$, where $T_{i}$’s are iid random variables from a (common) probability density function (PDF) $f_{\bm{\theta}}(t)$ and $\mathrm{I}(\cdot)$ is the event indicator function. Because all units within group $j$ share the same known entry time $\tau_{aj}$, but entry times differ across groups, the event indicator for group $j$ is evaluated relative to the group-specific time-on-test $\tau_{j}=t_{c}-\tau_{aj}$. The data observed at $t_{c}$ can be represented as a collection of single-cohort datasets, $\bm{D}=\{D_{1},\ldots,D_{g}\}$. For cohort $j$, we summarize the observed information as $D_{j}=(\bm{T}_{j},\delta_{j},\gamma_{j})$, where $\bm{T}_{j}$ is the vector of observed times possibly censored, $\delta_{j}$ denotes the number of failures, and $\gamma_{j}$ the number of units censored at $t_{c}$.
Let $Y$ be the number of item failures to be predicted at a future time $\tau_{j}+\Delta t$ on $n=\sum_{j=1}^{g}n_{j}$ tested unit. We denote $\zeta_{j}=\sum_{i=1}^{n_{j}}\mathrm{I}(\tau_{j}<T_{i}\leq\tau_{j}+\Delta t)$ as the indicator of the event occurrence between the (group) censoring time $\tau_{j}$ and the future time. Since $\zeta_{j}\sim\mathrm{Bin}\left(\pi_{j},\gamma_{j}\right)$ are mutually independent and $\pi_{j}=P(\tau_{j}<T_{j}\leq\tau_{j}+\Delta t\mid T_{j}>\tau_{j})$, then
@@
Following [33], Tian et al. [71] proposed obtaining the $1-\alpha$ prediction interval for $Y$ as the corresponding quantiles of the bootstrap estimator of the conditional CMF. Specifically, one generates $B$ bootstrap replications of the multiple-cohort data and, for each replication $b=1,\dots,B$, computes the vector $\hat{\boldsymbol{\pi}}^{*}_{b}$ of cohort-specific conditional event-probability. For each bootstrap sample, one then evaluates the CMF of $Y$, denotes as $F_{Y}(y\mid\hat{\boldsymbol{\pi}}^{*}_{b})$. Averaging these over $b$ yields the direct-bootstrap estimator of the conditional CMF is [71]
$\hat{F}^{*}_{Y}(y\mid\bm{D})=\frac{1}{B}\sum_{b=1}^{B}F_{Y}(y\mid\hat{\boldsymbol{\pi}}^{*}_{b}).$ (3)
@@ 3.1 Methods
At the time of the interim $t_{c}$ there are $n$ patient available (accrual closed) with different time origin of study entry $\mathcal{T}_{aj}$, with $j=1\ldots n$. Let $T_{j}$ denote the time to event, $L_{j}$ the time to lost to follow-up from the study and $\mathcal{T}_{j}=t_{c}-\mathcal{T}_{aj}$ the interim censoring time. Let $C_{j}=\mathrm{min}(L_{j},\mathcal{T}_{j})$ be the time to censoring, $T_{j}^{\mathrm{obs}}=\mathrm{min}(T_{j},C_{j})$ and $\delta_{j}$, $\varepsilon_{j}$, $\gamma_{j}$ be the indicator functions for $T_{j}\leq C_{j}$, $L_{j}\leq\mathrm{min}(T_{j},\mathcal{T}_{j})$, $\mathcal{T}_{j}\leq\mathrm{min}(T_{j},L_{j})$ respectively. Last, we denote $\bm{z}_{j}$ as the vector of patient-level covariates. We focus here on baseline charateristics.
We assume the follwing model for $T_{j}$, $L_{j}$ and $\mathcal{T}_{j}$:
$\begin{gathered}T_{j}\sim f_{\bm{\theta}}(t\mid\bm{z}_{j})\\[4.0pt] C_{j}=(L_{j},\mathcal{T}_{j})\sim g_{\bm{\psi},\bm{\phi}}(l,\tau)\\[4.0pt] (T_{j}\perp C_{j})\mid\bm{z}_{j}\\[4.0pt] \bm{\vartheta}=(\bm{\theta},\bm{\psi},\bm{\phi})\in\Theta\times\Psi\times\Phi\end{gathered}$ (5)
where $f_{\bm{\theta}}(t\mid\bm{z}_{j})$ denotes the PDF of the time to event, $g_{\bm{\psi},\bm{\phi}}(l,\tau)$ is the PDF of the censoring distribution expressed as the joint PDF of the time to lost to follow-up and the time of censoring at the interim. Since patients enter into the study at random times, the PDF of $\mathcal{T}_{j}$ is the conditional PDF of the entry times $\mathcal{T}_{aj}$ shifted by $t_{c}$ [43]. Thus, we assume an independent right-censoring mechanism, identical across treatment groups and covariate levels and that the PDFs of $T_{j}$, $L_{j}$ and $\mathcal{T}_{j}$ do not share any parameters. In contrast, interim censoring and lost to follow-up are not required to be independent processes, that is, we do not require that $g_{\bm{\psi},\bm{\phi}}(l,\tau)=\tilde{g}_{\bm{\psi}}(l)\,r_{\bm{\phi}}(\tau)$, where $g_{\bm{\psi}}(l)$ and $r_{\bm{\phi}}(\tau)$ are the marginal distribution of $L_{j}$ and $\mathcal{T}_{j}$ respectively.
Let $Y$ be the number of events to be predicted at a future time $\tau_{j}+\Delta t$ where $\tau_{j}=t_{c}-\tau_{aj}$ denotes a realization of $\mathcal{T}_{j}$. Furthermore, let $\zeta_{j}$ be the indicator function for $T_{j}\leq\mathrm{min}(\tau_{j}+\Delta t,L_{j})$. The probability that a patient will experience the event by the future time $\tau_{j}+\Delta t$, given that neither he/she has not yet experienced the event at the time of the interim $t_{c}$ neither was lost to follow-up is $\pi_{j}=\Pr\bigl(T_{j}\leq\mathrm{min}(\tau_{j}+\Delta t,L_{j})\mid\gamma_{j}=1,\bm{z}_{j}\bigr)$. Under assumptions (5), $\pi_{j}$ can be expressed as
@@
where $F_{\bm{\theta}}$ and $\tilde{G}_{\bm{\psi}}$ are the CFD of the event time and lost to follow-up time respectively. Proof of (6) and special cases are given in Appendix, Section  7.3. Since $\zeta_{j}\sim\mathrm{Ber}\left(\pi_{j}\right)=\mathrm{Bin}\left(\pi_{j},1\right)$ and are assumed mutually independent, the number of events to be predicted at a future time $\tau_{j}+\Delta t$ has still a Poisson-binomial distribution, that is
@@
were $\bm{{\pi}}=\{\pi_{j}\;:\;\gamma_{j}=1\}$. Prediction bounds are then obtained as in (4) after having recovered the bootstap estimator of the conditional CMF. Once $\bm{D}=(T_{j}^{\mathrm{obs}},\delta_{j},\varepsilon_{j},\gamma_{j})_{j=1}^{n}$ is observed, estimates for $\bm{\theta}$ and $\bm{\psi}$ are found under assumptions (5) by maximizing the following likelihoods
@@
Note that maximizing the likelihood in (8) is justified and yields the maximum likelihood estimator of $\bm{\psi}$. This implies that, for the sole purpose of construct prediction intervals in a within-sample prediction framework, we can ignore the dependence that may arise between the loss-to-follow-up time and the entry time when constructing prediction intervals, e.g., patients enrolled near the end of accrual may therefore be less likely to be lost to follow-up (Jiang, 2021). This justification is given in the Appendix, Section  7.4 and it is valid under assumptions (5).

### v2

@@ Abstract
Time-to-event endpoints are central to evaluate treatment efficacy across many disease areas. Many trial protocols include interim analyses within group-sequential designs that control type I error via spending functions or boundary methods, with operating characteristics determined by the number of looks and the information accrued. Planning interim analyses with time-to-event endpoints is challenging because statistical information depends on the number of observed events, so adequate follow-up to accrue the required events is critical and interim prediction of information at scheduled looks and at the final analysis becomes essential. While several methods have been developed to predict the calendar time required to reach a target number of events, to the best of our knowledge there is no established framework that addresses the prediction of the number of events at a future date with corresponding prediction intervals. Starting from prediction interval approach originally developed in reliability engineering for the number of future component failures, we reformulated and extended it to the context of interim monitoring in clinical trials. This adaptation yields a general framework for event-count prediction intervals in the clinical setting, taking the patient as the unit of analysis and accommodating a range of parametric survival models, patient-level covariates, stagged entry and possible dependence between entry dates and loss to follow-up. Prediction intervals are obtained in a frequentist framework from a bootstrap estimator of the conditional distribution of future events. The performance of the proposed approach is investigated via simulation studies and illustrated by analyzing a real-world phase III trial in childhood acute lymphoblastic leukaemia.
00footnotetext: $\dagger$ Correspondence: e.ratti9@campus.unimib.it
@@ 1 Introduction
Time-to-event endpoints are often used for evaluating treatment efficacy in many disease areas. However, designing trials in which the primary endpoint is a time-to-event variable is challenging. In this context, the statistical information of interest depends on the number of events that occur rather than on the number of patients enrolled. Accordingly, the trial is usually powered on the expected number of events [80]. Interim analyses are increasingly planned to allow early stopping for efficacy or futility for ethical and economic reasons [77]. Group-sequential designs are typically adopted and type-I error is controlled via alpha-spending or boundary methods [42]. The operating characteristics of a group sequential test - power, error rates, and stopping probabilities - depend on the number and timing of looks, on the information accrued at each analysis and on the maximum planned information at the final analysis. Consequently, the protocol must ensure sufficient follow‐up to achieve the target number of events required to preserve the nominal power. In addition, interim analyses should be conducted at the information levels specified at the design stage. If analyses are performed at times corresponding to a different number of events than planned, the Type I error probability may be adversely affected [46] [60].
@@
However, achieving both aims is difficult in practice. Sample size calculation relies on the correct specification of the survival distribution, yet trustworthy informations on survival estimates are often unavailable at the design stage and may shift as data mature. When the design assumptions prove inaccurate while the trial is ongoing, the operational characteristics of the study could be jeopardize. Since information accrues when events occur, an unforeseen slowdown in event rates can shift interim analyses to information times that depart from the protocol plan or push completion well beyond the planned timeline. This has an impact on the budget and logistics of the trial, even when the nominal power of the trial is preserved. On the other hand, an unexpected increase in event rates may put the trial at risk of being over-run [78] and appropriate methods should be considered [7]. A disciplined interim monitoring framework can identify such discrepancies early and support timely corrective actions.
@@
In light of these challenges, it is important to prospectively predict while the trial is ongoing the information expected at each planned interim analysis, the maximum (final) information at the scheduled final analysis or the calendar time at which these targets will be reached. If a maximum-information approach is adopted, the trial ends once the target number of events has accrued [44]. A natural predictand is therefore the calendar time required to reach that target, often termed milestone or time to endpoint maturation [57] [76]. In a maximum-duration approach, follow-up is limited by a fixed calendar time [44]. In this case, natural predictands are the cumulative number of events or, equivalently, the additional number of events expected by future calendar dates and at the planned study end.
@@
The point and interval prediction of the time to endpoint maturation has been extensively studied in the literature [6, 83, 81, 82, 66, 16, 47, 52]. Heitjan et al. [34] provide a comprehensive review of existing methods, highlighting the importance of prediction intervals for time to endpoint maturation to quantify uncertainty around the study’s closing date. Some approaches accommodate ongoing accrual at the time of prediction [76, 85, 3].
@@
Predicting the additional number of events at interim time points is critical for decision-making, especially when new drugs are being developed at a fast pace and adaptive and complex trials designs are increasingly needed, e.g. platform trials. Neverthells, most available methods in the biostatistical literature provide only point prediction under a few survival distributions. Fang and Su [24] proposed a hybrid approach that combines nonparametric and parametric modeling to estimate the survival curve. Specifically, change points in the survival function are first detected using a data-driven algorithm developed by Goodman et al. [28]. The segment preceding the final change point is estimated nonparametrically, while the tail beyond it is modeled using a parametric exponential or Weibull distribution. This approach is implemented by Rufibach [63] in the eventTrack R package. A growing body of literature has also addressed the use of point prediction for the purpose of sample size re-estimation [79, 31, 72, 25, 56]. Even though point predictions are easy to communicate they do not quantify uncertainty about reaching the planned information level and target number of events by the prespecified date. Prediction intervals address this limitation by explicitly reflecting uncertainty through their limits and width, making them more informative for interim trial monitoring. To the best of our knowledge, only two approaches address directly the issue of providing prediction interval for the additional number of events. Anisimov [3] derived closed-form prediction intervals for the number of future events in ongoing multicenter time-to-event trials, explicitly accounting for both newly recruited and at-risk patients under varied recruitment scenarios. The approach models center-level accrual via a Poisson–gamma process and event dynamics with continuous-time finite Markov chains under an exponential model. Anisimov et al. [2] extended the framework to incorporate a cure fraction, under both exponential and Weibull assumptions. However, these approaches construct prediction intervals using the so-called plug-in method [55], i.e., they treat the model parameters as if they were known and simply replace them with their point estimates when computing the conditional distribution of the additional number of events. This has been shown to yield intervals whose coverage can be far from the nominal level [71]. Motivated by a recently completed phase III trial in childhood acute lymphoblast leukemia (ALL), we propose here a general framework to construct prediction intervals for the additional number of events at interim analysis in fixed follow-up times clinical trials by adapting and extending a method originally developed in reliability engineering for predicting component failures [71]. We extended the approach to the clinical setting by treating the patient (rather than a component cohort) as the unit of analysis, accommodating possible dependence between entry dates and loss to follow-up and leveraging patient-level covariates available for instance in case of at unblinded interim looks.
@@
This paper is organized as follows: Section 2 introduces the notation, the prediction framework and briefly reviews the methods originally proposed by Tian et al. [71] for obtain prediction intervals; Section 3 outlines our proposed adaptation and extensions; Section 4, presents the results of a simulation study to evaluate the performance of the modified framework; Section 5 include an application to a phase III trial in childhood acute lymphoblast leukemia (ALL); Section 6 discusses the implications of our findings and highlights potential extensions.
@@ 2 Background
In this paper, we adapt and extend the direct bootstrap method in the multiple-chort scenario. Tian et al. [71] shown that the method has stronger empirical performance than the calibration bootstrap. Moreover, unlike the generalized pivotal-quantity bootstrap—whose construction is primarily well-defined under Type II censoring and only approximated under Type I censoring—the direct bootstrap can accommodates more general censoring mechanisms. We therefore focus on this approach and briefly describe it below.
@@ 2.1 Prediction Intervals Definition
[…]
@@ 2.2 The Direct-bootstrap, Multiple Chort Method (DB-MC)
At the prediction time $t_{c}$, there are $g$ cohorts (groups) of items. Cohort $j$ contains $n_{j}$ items that were put on test at the common known entry time $\tau_{aj}$, with $j=1,\ldots,g$. Let $T_{ij}$ denote the failure time of item $i$ in cohort $j$, and assume
$T_{ij}\ \stackrel{{\scriptstyle\text{iid}}}{{\sim}}\ f_{\boldsymbol{\theta}}(t),\qquad i=1,\ldots,n_{j},\ \ j=1,\ldots,g,$ (2)
where all items share the same (common) PDF $f_{\boldsymbol{\theta}}(t)$. Define the event indicator at $t_{c}$ for item $i$ in cohort $j$ as $\tilde{\delta}_{ij}=\mathbb{I}\!\left(T_{ij}\leq\tau_{j}\right)$ where $\tau_{j}=t_{c}-\tau_{aj}$. The number of observed failures in cohort $j$ is $\delta_{j}=\sum_{i=1}^{n_{j}}\tilde{\delta}_{ij}$ and the number of right-censored items is $\gamma_{j}=n_{j}-\delta_{j}$. The data observed at $t_{c}$ can be represented as a collection of single-cohort datasets, $\bm{D}=\{D_{1},\ldots,D_{g}\}$. For cohort $j$, we summarize the observed information as $D_{j}=(\bm{T}_{j},\delta_{j},\gamma_{j})$, where $\bm{T}_{j}$ is the vector of observed times possibly right-censored, $\delta_{j}$ denotes the number of failures, and $\gamma_{j}$ the number of units censored at $t_{c}$.
Let $Y$ be the number of item failures to be predicted at a future time $\tau_{j}+\Delta t$ on $n=\sum_{j=1}^{g}n_{j}$ tested unit. We denote $\zeta_{j}=\sum_{i=1}^{n_{j}}\mathbb{I}(\tau_{j}<T_{ij}\leq\tau_{j}+\Delta t)$. Since $\zeta_{j}\sim\mathrm{Bin}\left(\pi_{j},\gamma_{j}\right)$ are mutually independent and $\pi_{j}=P(T_{j}\leq\tau_{j}+\Delta t\mid T_{j}>\tau_{j})$, then
@@
Following [33], Tian et al. [71] proposed obtaining the $1-\alpha$ prediction interval for $Y$ as the corresponding quantiles of the bootstrap estimator of the conditional CMF. Specifically, once $\bm{D}=\{D_{1},\ldots,D_{g}\}$ is observed, estimate $\bm{\hat{\theta}}$ for $\bm{\theta}$ is found by maximizing the parametric likelihood
$\mathcal{L}(\bm{\theta})$ $\propto\prod_{j=1}^{g}\prod_{i=1}^{n_{j}}f_{\bm{\theta}}(t)^{\tilde{\delta_{i}}}\Big[1-F_{\bm{\theta}}(t)\Big]^{1-\tilde{\delta_{i}}}$, (4)
where $F_{\bm{\theta}}(t)$ is the CDF of $T_{ij}$. Thus, one generates $B$ bootstrap replications of the multiple-cohort data from $f_{\bm{\hat{\theta}}}(t)$ and, for each replication $b=1,\dots,B$, computes the vector $\hat{\boldsymbol{\pi}}^{*}_{b}$ of cohort-specific conditional event-probability at a future prediction horizon $\Delta t$. For each bootstrap sample, one then evaluates the CMF of $Y$, denotes as $F_{Y}(y\mid\hat{\boldsymbol{\pi}}^{*}_{b})$. Averaging these over $b$ yields the direct-bootstrap estimator of the conditional CMF is [71]
$\hat{F}^{*}_{Y}(y\mid\bm{D})=\frac{1}{B}\sum_{b=1}^{B}F_{Y}(y\mid\bm{D}_{b}^{*};\hat{\boldsymbol{\pi}}^{*}_{b}).$ (5)
@@ 3.1 Methods
At the interim time $t_{c}$, accrual is closed and $n$ patients are available, each with a specific entry time $\mathcal{T}_{aj}$, $j=1,\ldots,n$. Let $T_{j}$ denote the time to event and $L_{j}$ the time to loss to follow-up, both measured from study entry. Define the individual censoring time at the interim as $\mathcal{T}_{j}=t_{c}-\mathcal{T}_{aj}$. Let $C_{j}=\min(L_{j},\mathcal{T}_{j})$, $T_{j}^{\mathrm{obs}}=\min(T_{j},C_{j}),$ and define the event-type indicators $\delta_{j}=\mathbb{I}(T_{j}\leq C_{j})$, $\varepsilon_{j}=\mathbb{I}\big(L_{j}\leq\min(T_{j},\mathcal{T}_{j})\big)$, $\gamma_{j}=\mathbb{I}\big(\mathcal{T}_{j}\leq\min(T_{j},L_{j})\big)$. Finally, let $\bm{z}_{j}$ denote the vector of patient-level covariates. Here we focus on baseline characteristics. We impose the following assumptions on $(T_{j},L_{j},\mathcal{T}_{j})$:
$\text{(A1) }\,\,T_{j}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\ f_{\bm{\theta}}(\,t\mid\bm{z}_{j}).$ (7)
$\text{(A2) }\,\,(L_{j},\mathcal{T}_{j})\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\ g_{\bm{\psi},\bm{\phi}}(l,\tau).$
$\text{(A3) }\,\,T_{j}\perp(L_{j},\mathcal{T}_{j})\ \mid\ \bm{z}_{j}.$
$\text{(A4) }\,\,\bm{\vartheta}=(\bm{\theta},\bm{\psi},\bm{\phi})\in\Theta\times\Psi\times\Phi.$
Note that, since patients enter the study at random calendar times, the distribution of $\mathcal{T}_{j}$ is induced by the distribution of entry times $\mathcal{T}_{aj}$, shifted by $t_{c}$ (i.e., $\mathcal{T}_{j}=t_{c}-\mathcal{T}_{aj}$); see [43]. Thus, we assume an independent right-censoring mechanism, identical across treatment groups and covariate levels and that the PDFs of $T_{j}$, $L_{j}$ and $\mathcal{T}_{j}$ do not share any parameters. In contrast, interim censoring and loss to follow-up are not required to be independent processes, that is, we do not require that $g_{\bm{\psi},\bm{\phi}}(l,\tau)=\tilde{g}_{\bm{\psi}}(l)\,r_{\bm{\phi}}(\tau)$, where $\tilde{g}_{\bm{\psi}}(l)$ and $r_{\bm{\phi}}(\tau)$ are the marginal distribution of $L_{j}$ and $\mathcal{T}_{j}$ respectively. This implies that we do not account for any potential dependence between the loss-to-follow-up times and the entry times, e.g., patients enrolled near the end of accrual may therefore be less likely to be loss to follow-up [43].
Let $Y$ be the number of events to be predicted at a future time $\tau_{j}+\Delta t$ where $\tau_{j}=t_{c}-\tau_{aj}$ denotes a realization of $\mathcal{T}_{j}$. Furthermore, let $\zeta_{j}$ be the indicator function for $T_{j}\leq\mathrm{min}(\tau_{j}+\Delta t,L_{j})$. The probability that a patient will experience the event by the future time $\tau_{j}+\Delta t$, given that neither he/she has not yet experienced the event at the time of the interim $t_{c}$ neither was loss to follow-up is $\pi_{j}=\Pr\bigl(T_{j}\leq\mathrm{min}(\tau_{j}+\Delta t,L_{j})\mid\gamma_{j}=1,\bm{z}_{j}\bigr)$. Under assumptions (7), $\pi_{j}$ can be expressed as
@@
where $F_{\bm{\theta}}$ and $\tilde{G}_{\bm{\psi}}$ are the CDFs of the event time and loss to follow-up time respectively. Proof of (8) and special cases are given in Appendix, Section  7.3. Since $\zeta_{j}\sim\mathrm{Ber}\left(\pi_{j}\right)=\mathrm{Bin}\left(\pi_{j},1\right)$ and are assumed mutually independent, the number of events to be predicted at a future time $\tau_{j}+\Delta t$ has still a Poisson-binomial distribution, that is
@@
were $\bm{{\pi}}$ is the vector of individual conditional probabilities obtained for only the interim censored patients as in (8). Prediction bounds are then obtained as in (6) after having recovered the bootstap estimator of the conditional CMF. Once the interim data $\bm{D}=(T_{j}^{\mathrm{obs}},\delta_{j},\varepsilon_{j},\gamma_{j},\bm{z}_{j})_{j=1}^{n}$ is observed, estimates for $\bm{\theta}$ and $\bm{\psi}$ are found by maximizing the following parametric likelihoods
@@
Maximizing the likelihoods in (10) is justified under assumptions (7) and yields the maximum likelihood estimator for $\bm{\theta}$ and $\bm{\psi}$. This justification is given in the Appendix, Section  7.4.